Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y).
3x + 2y = 2
-2x + y = 8

Answer by Alan3354(69443)

3x + 2y = 2
-2x + y = 8 --> y = 2x + 8
==================
Sub for y in the 1st equation.